A large sport utility vehicle has a mass of 2.5 x 10^3 kg. calculate the mass of CO2 emitted into the atmosphere upon acceleration the SUV from 0mph to 65.0 mph. assume that the required energy comes from the combustion of octane with 30% efficiency.
1
Expert's answer
2017-12-14T04:25:08-0500
m= 2,5*103 kg
η=30%
v0=65mil/hour= 29,06m/s
qC8H18 = 47,89*106 J/kg
η=A/Q *100%
A=F*t
Q= qC8H18*m C8H18
F=m*a a= (vk - v0)/t
F=m*(vk - v0)
A= m*vk
η= (m*vk / qC8H18*m C8H18)*100%
m C8H18= (m*vk/ η* qC8H18)*100% = (2,5*103 kg*29,06m/s)/(0,3*47,89*106 J/kg)=5,06*10-3kg
2C8H18+25O2=16CO2+18H2O
n (C8H18) = 5,06g/ 114g/mol =0,044mol
n(CO2) = 8*n (C8H18) = 0,355 mol
m(CO2) = n(CO2) *M =0,355 mol*44 g /mol = 15,62g